OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95

OpenStudy (anonymous):

i think 197

OpenStudy (anonymous):

197

OpenStudy (angela210793):

a18=a1+(n-1)*d 95=-7+17d 17d=102 d=6 a35=a1+(n-1)*d=-7+34*6=197

OpenStudy (angela210793):

Arithmetic progression: a2-a1=d an=a1+(n-1)*d Sn=(a1+an)*n/2 these r the formulas please try to solve the problems by urself 1st...

OpenStudy (anonymous):

i just told you that i do, and then i check them because im taking quizes and i wanna make sure that i have the right answer.

OpenStudy (angela210793):

ok...didn't know tht sorry

OpenStudy (anonymous):

its okay i appreciate everyones help you guys are awesome !